OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sine of pi divided by two times cosine of pi divided by seven plus cosine of pi divided by two times sine of pi divided by seven.

OpenStudy (solomonzelman):

write your equation out.

OpenStudy (anonymous):

okay give me a second please

OpenStudy (anonymous):

okay give me a second please \[\sin (\pi/2) \cos(\pi/7) + \cos(\pi/2) \sin(\pi/7)\]

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm sin(π/2+π/7) }\) this is your expression :)

OpenStudy (solomonzelman):

Now, simplify it.

OpenStudy (neer2890):

sin(x+y)= sinx cosy+cosx siny

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm sin(π/2+π/7) }\) \(\large\color{midnightblue}{ \rm sin(7π/14+2π/14) }\) \(\large\color{midnightblue}{ \rm sin(9π/14) }\)

OpenStudy (anonymous):

So is it sin(9pi/14)?

OpenStudy (anonymous):

Oh okay yes thanks so much @SolomonZelman and @neer2890

OpenStudy (solomonzelman):

anytime